[figure: 9.7 cm]

INFLUENCE OF FILM THICKNESS ON THE MICROSTRUCTURE AND MAGNETIC PROPERTIES OF FINEMETIC THIN FILMS

R. Dlugoš^{1)}, P. Sovák^{1)}, J. Balcerski^{2)}, J. Füzer^{1)}

1)Faculty of Science, P.J.Šafárik University, Park Angelinum 9, 041 54 Košice, Slovakia 2)Department of Solid State Physics University of Łódź, Pomorska 141, Łódź, Poland

The aim of this work was to study the influence of film thickness on the structure and magnetic properties of finemetic thin films after annealing. Thin films with various thickness (from 20nm up to 1000nm) were prepared by DC sputtering method. Heat treatments of the films for further structural and magnetic observations were performed at the temperature range 300-550^{o}C for 15 minutes in a vacuum furnace. Structural observations were carried out by X-ray diffractometry and transmission electron microscopy. The bulk value of coercivity was determined from hysteresis loops traced with fluxmeter. All the experimental results confirm a different magnetic behaviour of the thin films according to their thickness.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : oral or poster

Corresponding author : Rastislav Dlugoš

Address for correspondence : Department of Experimental Physics Faculty of Sciences P.J.Šafárik University Park Angelinum 9 041 54 Košice Slovak Republic

Email address : dlugos@kosice.upjs.sk